# SalUn: Empowering Machine Unlearning via Gradient-based Weight Saliency in Both Image Classification and Generation

### 19 Oct 2023 v1

@@ 3 Preliminaries and Problem Statement
To be concrete, let $\mathcal{D}=\{\mathbf{z}_{i}\}_{i=1}^{N}$ denote the training dataset encompassing $N$ data points (including data feature $\mathbf{x}_{i}$ and label $\mathbf{y}_{i}$ for supervised learning). And let $\mathcal{D_{\mathrm{f}}}\subseteq\mathcal{D}$ be the forgetting dataset. Its complement, denoted by $\mathcal{D_{\mathrm{r}}}=\mathcal{D}\setminus\mathcal{D}_{\mathrm{f}}$, is referred to as the remaining dataset. Prior to MU, we denote the original model as ${\bm{\theta}}_{\mathrm{o}}$}, trained on $\mathcal{D}$ using, e.g., empirical risk minimization (ERM). Consistent with existing literature (Thudi et al., 2022a), we regard Retrain as the MU gold standard, which involves training mode parameters ($\bm{\theta}$}) from scratch over $\mathcal{D}_{\mathrm{r}}$. Nonetheless, Retrain can be computationally demanding. Hence, the central challenge in MU is to acquire an unlearned model (referred to as ${\bm{\theta}}_{\mathrm{u}}$}) from ${\bm{\theta}}_{\mathrm{o}}$} on $\mathcal{D}_{\mathrm{f}}$ and/or $\mathcal{D_{\mathrm{r}}}$, so that it can serve as an accurate and computationally efficient substitute for Retrain. In what follows, we introduce two MU paradigms that are the primary focus of this work: MU for image classification and MU for image generation.
@@
Evaluating the effectiveness of MU for image classification has involved the use of various metrics. While a consensus is still lacking, we adhere to the recent approach proposed by (Jia et al., 2023), which considers a comprehensive ‘full-stack’ MU evaluation. This includes unlearning accuracy (UA), i.e., the accuracy measure of an unlearned model ${\bm{\theta}}_{\mathrm{u}}$} on $\mathcal{D}_{\mathrm{f}}$, membership inference attack (MIA) on $\mathcal{D}_{\mathrm{f}}$, i.e., the privacy measure of ${\bm{\theta}}_{\mathrm{u}}$} over $\mathcal{D}_{\mathrm{f}}$, remaining accuracy (RA), i.e., the fidelity of an unlearned model ${\bm{\theta}}_{\mathrm{u}}$} on the remaining training set $\mathcal{D}_{\mathrm{r}}$, testing accuracy (TA), i.e., the generalization of ${\bm{\theta}}_{\mathrm{u}}$}, and run-time efficiency (RTE), i.e., the computation time of applying an MU method.
@@ 4 Challenges in Current Machine Unlearning Methods
Figure 2: Demonstration of the instability limitations of MU methods on CIFAR-10. (a) Sensitivity of unlearning accuracy gaps with respect to Retrain (measured by $|\text{Method}-\text{{\text{Retrain}}}|$’) as a function of forgetting data amount. Four MU methods (FT, RL, IU, $\ell_{1}$-sparse) are included. (b) Box plots illustrating the variances of unlearning accuracies using Retrain, IU, and the proposed weight saliency-integrated IU across various hyperparameter choices. The box size represents the variance of UA values against hyperparameter values.
@@
The instability limitation. In evaluating the performance of MU methods, previous research has often assumed a fixed number of forgetting data, such as data points within an entire class or a fixed ratio of the training set. There has been limited evaluation exploring how the unlearning performance is affected by varying quantities of forgetting data. In Fig. 2(a), we investigate the unlearning performance gap relative to the gold standard Retrain, measured in terms of UA (unlearning accuracy), as a function of the quantity of forgetting data points. Note that a smaller performance gap is desirable. As we can see, the unlearning effectiveness of MU methods (①-⑤) observed at a 10% forgetting data quantity does not necessarily hold when the forgetting data quantity is increased to 50%. Similar instability in other performance metrics can also be observed in more experiment results later.
@@ 5 SalUn: Weight Saliency Is Possibly All You Need for MU
Gradient-based weight saliency map. We first illustrate the rationale behind exploring gradient-based weight saliency for MU. Recent evidence suggests that ML models exhibit modularity characteristics to some extent (Menik & Ramaswamy, 2023). Here modularity refers to the property of a large ML model being decomposable into manageable subparts, each of which can be more easily maintained and updated independently. In particular, weight sparsity (Frankle & Carbin, 2018) has gained recognition as an important driver of modularity, leading to improvements in various aspects of ML, including efficiency (Riquelme et al., 2021), interpretability (Wong et al., 2021), and robustness (Chen et al., 2022b). In the context of MU, weight sparsity has also been harnessed to facilitate the unlearning process, leading to the $\ell_{1}$-sparse unlearning baseline (Jia et al., 2023). However, weight sparsity encounters certain limitations when applied to MU: (1) Determining the appropriate sparse pattern for an ML model (e.g., a DM) can be a challenging task in itself; (2) Even when sparsity is achievable, some applications may not favor delivering a sparse model after MU due to the observed performance decline, as exemplified by the $\ell_{1}$-sparse MU method in Sec. 4.
@@
As shown in(3) and (4), weight saliency offers a principled method to tackle the optimization problem in MU when the forgetting loss $\ell_{\mathrm{f}}({\bm{\theta}};\mathcal{D}_{\mathrm{f}})$ is specified. It is worth noting that the forgetting loss $\ell_{\mathrm{f}}$ used in existing MU methods can be considered a suitable candidate for calculating the weight saliency map (3). In this study, we find that the forgetting loss in GA (gradient ascent) (Thudi et al., 2022a) presents a unified, effective, and simple solution in image classification and generation:
@@
where $\ell_{\mathrm{CE}}$ is the cross-entropy (CE) loss for supervised classification, and $\ell_{\mathrm{MSE}}$ has been defined in DM training (2). The weight saliency map for MU can be then obtained through (3) and (5).
Saliency-based unlearning (SalUn). Next, we introduce SalUn, which incorporates (4) into the unlearning process. One advantage of SalUn is its plug-and-play capability, allowing it to be applied on top of existing unlearning methods. In particular, we find that integrating weight saliency with the RL (random labeling) method provides a promising MU solution; See an ablation study in Table. A1.
@@
Extension on ‘soft-thresholding’ SalUn. The implementation of SalUn relies on a pre-selected hard threshold to determine the weight saliency map $\mathbf{m}_{\mathrm{S}}$ in (3). While this hard-thresholding approach performs well, we can also develop an alternative implementation of SalUn that employs soft thresholding and may allow for a more flexible saliency map determination; See Appendix A for more algorithmic details. Yet, in practice, the soft-thresholding variant does not outperform the hard-thresholding version. Thus, we will focus on the hard-thresholding implementation by default.
@@ 6.1 Experiment setups
Table 1: Performance summary of various MU methods for image classification (including the proposed SalUn and SalUn-soft and 7 other baselines) in two unlearning scenarios, 10% random data forgetting and 50% random data forgetting, on CIFAR-10 using ResNet-18. The result format is given by $a_{\pm b}$, with mean $a$ and standard deviation $b$ over 1010 independent trials. A performance gap against Retrain is provided in ($\bullet$). Note that the better performance of an MU method corresponds to the smaller performance gap with Retrain. The metric averaging (avg.) gap is introduced and calculated by the average of the performance gaps measured in accuracy-related metrics, including UA, MIA, RA, and TA. RTE is in minutes. More results of random data forgetting on CIFAR-10 using ResNet-18 are listed in Table A4.
@@
Baselines and evaluation. In our experiments, we will cover 9 unlearning baselines, which include 5 existing baselines presented in Sec. 4, FT (Warnecke et al., 2021), RL (Golatkar et al., 2020), GA (Thudi et al., 2022a), IU (Izzo et al., 2021; Jia et al., 2023), $\ell_{1}$-sparse (Jia et al., 2023), as well as 4 new baselines, including 2 boundary unlearning methods (Chen et al., 2023), boundary shrink (BS) and boundary expanding (BE), and 2 concept-unlearning methods, erased stable diffusion (ESD) (Gandikota et al., 2023) and forget-me-not (FMN) (Zhang et al., 2023). In our performance evaluation for image classification, we adhere to the 5 evaluation metrics described in Sec. 3. We use UA and MIA to measure the unlearning efficacy, RA and TA to assess the unlearned classifier’s fidelity and generalization ability, and RTE to evaluate the computation efficiency in MU. In the context of image generation, we train an external classifier to evaluate UA, ensuring that the generated images do not belong to the forgetting class/concept. We adopt ResNet-34 trained on CIFAR-10 and a pre-trained ResNet-50 on ImageNet. Besides UA on the forgetting data, we also use FID to evaluate the quality of image generations for non-forgetting classes/prompts.
@@ 6.2 Experiment results
Performance of MU in image classification. In Table 1, we present a comprehensive comparison between our proposed method (SalUn and its soft-thresholding variant referred to as ‘SalUn-soft’) and 7 other MU baselines (FT, RL, GA, IU, $\ell_{1}$-sparse, BS, and BE) designed for image classification. Motivated by the instability limitation at the rise of the quantity of forgetting data points as discussed in Sec. 4, we explore two unlearning cases: the standard $10\%$ random data forgetting and the higher $50\%$ random data forgetting. The unlearning performance is evaluated using the five metrics introduced earlier: UA, MIA, RA, TA, and RTE. Moreover, we include the performance of the exact unlearning method Retrain for comparison. It’s important to note that a better approximate unlearning method should exhibit a smaller performance gap compared to Retrain. To quantify the reduction in the performance gap, we introduce the metric called the averaging (avg.) gap, calculated as the average of the performance gaps measured in accuracy-related metrics, including UA, MIA, RA, and TA. We draw some key observations from Table 1 below.
@@
Figure 4: Image generations of using SalUn and its random weight saliency masking variant (we call ‘random’) for DDPM on CIFAR-10. The forgetting class is given by ‘airplane’, ‘I’ refers to the generated image sample under the class condition ‘airplane’, and ‘C’ refers to the non-forgetting class name, e.g. ‘car’ (C1).
@@
Weight saliency mask is key to adapting MU for image generation. In Fig. 4, we explore the influence of the weight saliency mask in SalUn when transitioning to image generation. For comparison, we also include the random masking baseline. As we can see, the use of random masking in SalUn could lead to unstable generation performance when the condition is set to the forgetting class ‘airplane’ or other non-forgetting classes. First, the generation of noisy images may indicate over-forgetting, which contradicts the results obtained with Retrain as shown in Fig. 3. Second, when a non-forgetting class is specified, we notice the use of random masking degrades the generation quality, e.g., $C2$, $C3$, $C6$, and $C7$ in the figure. In contrast, our proposed SalUn, which leverages a proper weight saliency map, outperforms the implementation that uses random masking. This highlights the significance of proper weight saliency in MU for image generation.
Extended from Fig. 4, we quantify the unlearning performance of Retrain, ESD, and SalUn using the two metrics introduced earlier, FID and UA, in Appendix B.3. With a comparable UA performance, it is worth highlighting that SalUn significantly outperforms ESD in FID. ESD seems to exhibit instability when forgetting and learning low-quality images like CIFAR-10.
Figure 5: Effectiveness of ‘nudity’ removal using different unlearned SD models acquired by SalUn, ESD, and FMN, respectively, and the original SD V1.4. The performance is measured by the # of generated harmful images against I2P prompts within each nudity category (i.e., row name).
@@
Application to NSFW concept forgetting. Further, we assess the effectiveness of SalUn in concept-wise forgetting to eliminate the impact of NSFW (not safe for work) concepts introduced through inappropriate image prompts (I2P) (Schramowski et al., 2023). Specifically, we generate images from the open-source SD V1.4 using prompts provided by I2P and classify them into various corresponding nude body parts using the NudeNet detector (Bedapudi, 2019). Our goal is to leverage an MU method to effectively erase the influence of the nudity-related prompts in SD. Fig. 5 presents the unlearning performance of different unlearning methods, including SalUn, and the ESD and FMT baselines introduced in Table 2. Here, the unlearning effectiveness is measured by the amount of nudity-related image generations using the unlearned SD with I2P prompts. We also present the performance of the original SD model for comparison. As we can see, SalUn generates the fewest harmful images across all the nude body part classes. In particular, it significantly outperforms ESD (the second best-performing method) in ‘male breast’ and ‘female breast’. Additionally, in the absence of unlearning, the original SD V1.4 can generate a large number of harmful images, underscoring the importance of MU for image generation. Fig. 6 provides a set of generated images using I2P prompts, illustrating the generation differences among the original SD model and its various unlearned versions.
@@
Figure 6: Examples of generated images using SDs w/ and w/o MU. The unlearning methods include ESD, FMN, and SalUn (ours). Each column represents generated images using different SDs with the same prompt (denoted by PPi) and the same seed. The specific descriptions of the prompts used are provided in Table A9.

### 06 Nov 2023 v2

@@ 3 Preliminaries and Problem Statement
To be concrete, let $\mathcal{D}=\{\mathbf{z}_{i}\}_{i=1}^{N}$ denote the training dataset encompassing $N$ data points (including data feature $\mathbf{x}_{i}$ and label $\mathbf{y}_{i}$ for supervised learning). And let $\mathcal{D_{\mathrm{f}}}\subseteq\mathcal{D}$ be the forgetting dataset. Its complement, denoted by $\mathcal{D_{\mathrm{r}}}=\mathcal{D}\setminus\mathcal{D}_{\mathrm{f}}$, is referred to as the remaining dataset. Prior to MU, we denote the original model as ${\bm{\theta}}_{\mathrm{o}}$}, trained on $\mathcal{D}$ using, e.g., empirical risk minimization (ERM). Consistent with existing literature (Thudi et al., 2022a), we regard Retrain as the MU gold standard, which involves training model parameters ($\bm{\theta}$}) from scratch over $\mathcal{D}_{\mathrm{r}}$. Nonetheless, Retrain can be computationally demanding. Hence, the central challenge in MU is to acquire an unlearned model (referred to as ${\bm{\theta}}_{\mathrm{u}}$}) from ${\bm{\theta}}_{\mathrm{o}}$} on $\mathcal{D}_{\mathrm{f}}$ and/or $\mathcal{D_{\mathrm{r}}}$, so that it can serve as an accurate and computationally efficient substitute for Retrain. In what follows, we introduce two MU paradigms that are the primary focus of this work: MU for image classification and MU for image generation.
@@
Evaluating the effectiveness of MU for image classification has involved the use of various metrics. While a consensus is still lacking, we adhere to the recent approach proposed by (Jia et al., 2023), which considers a comprehensive ‘full-stack’ MU evaluation. This includes unlearning accuracy (UA), (i.e., 1 - accuracy of an unlearned model ${\bm{\theta}}_{\mathrm{u}}$} on $\mathcal{D}_{\mathrm{f}}$), membership inference attack (MIA) on $\mathcal{D}_{\mathrm{f}}$, i.e., the accuracy of leveraging MIA to determine if a forgetting data is not in the training set of ${\bm{\theta}}_{\mathrm{u}}$}, remaining accuracy (RA), i.e., the fidelity of an unlearned model ${\bm{\theta}}_{\mathrm{u}}$} on the remaining training set $\mathcal{D}_{\mathrm{r}}$, testing accuracy (TA), i.e., the generalization of ${\bm{\theta}}_{\mathrm{u}}$}, and run-time efficiency (RTE), i.e., the computation time of applying an MU method. Except for RTE, a good unlearning method is expected to achieve performance close to that of Retrain.
@@ 4 Challenges in Current Machine Unlearning Methods
Figure 2: Demonstration of the instability limitations of MU methods on CIFAR-10. (a) Sensitivity of unlearning accuracy gaps with respect to Retrain (measured by $|\text{Method}-\text{{\text{Retrain}}}|$’) as a function of forgetting data amount. Five existing MU methods (FT, RL, GA, IU, $\ell_{1}$-sparse) are included. (b) Box plots illustrating the variances of unlearning accuracies using Retrain, IU, and the proposed weight saliency-integrated IU across various hyperparameter choices. The box size represents the variance of UA values across hyperparameter values.
@@
The instability limitation. In evaluating the performance of MU methods, previous research has often assumed a fixed number of forgetting data, such as data points within an entire class or a fixed fraction of the training set. There has been limited evaluation exploring how the unlearning performance is affected by varying quantities of forgetting data. In Fig. 2(a), we investigate the unlearning performance gap relative to the gold standard Retrain, measured in terms of UA (unlearning accuracy), as a function of the quantity of forgetting data points. Note that a smaller performance gap is desirable. As we can see, the unlearning effectiveness of MU methods (①-⑤) observed at a 10% forgetting data quantity does not necessarily hold when the forgetting data quantity is increased to 50%. Similar instability in other performance metrics can also be observed in more experiment results later.
@@ 5 SalUn: Weight Saliency Is Possibly All You Need for MU
Gradient-based weight saliency map. We first illustrate the rationale behind exploring gradient-based weight saliency for MU. Recent evidence suggests that ML models exhibit modularity characteristics to some extent (Menik & Ramaswamy, 2023). Here modularity refers to the property of a large ML model being decomposable into manageable subparts, each of which can be more easily maintained and updated independently. In particular, weight sparsity (Frankle & Carbin, 2018) has gained recognition as an important driver of modularity, leading to improvements in various aspects of ML, including efficiency (Riquelme et al., 2021), interpretability (Wong et al., 2021), and robustness (Chen et al., 2022b). In the context of MU, weight sparsity has also been harnessed to facilitate the unlearning process, leading to the $\ell_{1}$-sparse unlearning baseline (Jia et al., 2023). However, weight sparsity encounters certain limitations when applied to image generation: (1) Determining the appropriate sparse pattern for an ML model (e.g., a DM) can be a challenging task in itself; (2) Even when sparsity is achievable, some applications may not favor delivering a sparse model after MU due to the observed performance decline, as exemplified by the $\ell_{1}$-sparse MU method in Sec. 4.
@@
As shown in(3) and (4), weight saliency offers a principled method to tackle the optimization problem in MU when the forgetting loss $\ell_{\mathrm{f}}({\bm{\theta}};\mathcal{D}_{\mathrm{f}})$ is specified. It is worth noting that the forgetting loss $\ell_{\mathrm{f}}$ used in existing MU methods can be considered a suitable candidate for calculating the weight saliency map (3). In this study, we find that the forgetting loss in GA (gradient ascent) (Thudi et al., 2022a) presents a unified, effective, and simple solution in image classification and generation. This is given by the training loss over the forgetting dataset $\mathcal{D}_{\mathrm{f}}$:
@@
where $\ell_{\mathrm{CE}}$ is the cross-entropy (CE) loss for supervised classification, and $\ell_{\mathrm{MSE}}$ has been defined in DM training (2). The weight saliency map for MU can be then obtained by applying (5) to (3).
Saliency-based unlearning (SalUn). Next, we introduce SalUn, which incorporates the saliency-aware unlearning variables ${\bm{\theta}}_{\mathrm{u}}$ (4) into the unlearning process. One advantage of SalUn is its plug-and-play capability, allowing it to be applied on top of existing unlearning methods. In particular, we find that integrating weight saliency with the RL (random labeling) method provides a promising MU solution; See an ablation study in Table. A1.
@@
Extension on ‘soft-thresholding’ SalUn. The implementation of SalUn relies on a pre-selected hard threshold to determine the weight saliency map $\mathbf{m}_{\mathrm{S}}$ in (3). While this hard-thresholding approach performs well in practice, we can also develop an alternative implementation of SalUn that employs soft thresholding and could allow for a more flexible saliency map determination; See Appendix A for more algorithmic details. Yet, in practice, the soft-thresholding variant does not outperform the hard-thresholding version. Thus, we will focus on the hard-thresholding implementation by default.
@@ 6.1 Experiment setups
Table 1: Performance summary of various MU methods for image classification (including the proposed SalUn and SalUn-soft and 7 other baselines) in two unlearning scenarios, 10% random data forgetting and 50% random data forgetting, on CIFAR-10 using ResNet-18. The result format is given by $a_{\pm b}$, with mean $a$ and standard deviation $b$ over $10$ independent trials. A performance gap against Retrain is provided in ($\bullet$). Note that the better performance of an MU method corresponds to the smaller performance gap with Retrain. The metric averaging (avg.) gap is introduced and calculated by the average of the performance gaps measured in accuracy-related metrics, including UA, MIA, RA, and TA. RTE is in minutes. More results of random data forgetting on CIFAR-10 using ResNet-18 are listed in Table A4.
@@
Baselines and evaluation. In our experiments, we will cover 9 unlearning baselines, which include 5 existing baselines presented in Sec. 4, FT (Warnecke et al., 2021), RL (Golatkar et al., 2020), GA (Thudi et al., 2022a), IU (Izzo et al., 2021; Jia et al., 2023), $\ell_{1}$-sparse (Jia et al., 2023), as well as 4 new baselines, including 2 boundary unlearning methods (Chen et al., 2023), boundary shrink (BS) and boundary expanding (BE), and 2 concept-unlearning methods, erased stable diffusion (ESD) (Gandikota et al., 2023) and forget-me-not (FMN) (Zhang et al., 2023). In our performance evaluation for image classification, we adhere to the 5 evaluation metrics described in Sec. 3. We use UA and MIA to measure the unlearning efficacy, RA and TA to assess the unlearned classifier’s fidelity and generalization ability, and RTE to evaluate the computation efficiency in MU. In the context of image generation, we train an external classifier to evaluate UA, ensuring that the generated images do not belong to the forgetting class/concept. We adopt ResNet-34 trained on CIFAR-10 and a pre-trained ResNet-50 on ImageNet. Besides UA on the forgetting data, we also use FID (Fréchet inception distance) to evaluate the quality of image generations for non-forgetting classes/prompts.
@@ 6.2 Experiment results
Performance of MU in image classification. In Table 1, we present a comprehensive comparison between our proposed method (SalUn and its soft-thresholding variant referred to as ‘SalUn-soft’) and 7 other MU baselines (FT, RL, GA, IU, $\ell_{1}$-sparse, BS, and BE) designed for image classification. Motivated by the instability limitation with respect to the quantity of forgetting data points as discussed in Sec. 4, we explore two unlearning cases: the standard $10\%$ random data forgetting and the higher $50\%$ random data forgetting. The unlearning performance is evaluated using the five metrics introduced earlier: UA, MIA, RA, TA, and RTE. Moreover, we include the performance of the exact unlearning method Retrain for comparison. It’s important to note that a better approximate unlearning method should exhibit a smaller performance gap compared to Retrain. To quantify the reduction in the performance gap, we introduce the metric called the averaging (avg.) gap, calculated as the average of the performance gaps measured in accuracy-related metrics, including UA, MIA, RA, and TA. We draw some key observations from Table 1 below.
@@
Figure 4: Image generations of SalUn and a random masking baseline (‘random’) for DDPM on CIFAR-10. The forgetting class is given by ‘airplane’, ‘I1’, …, ‘I4’ refer to generated image samples under the class condition ‘airplane’, and ‘C1’, …, ‘C9’ refer to non-forgetting classes, e.g. ‘car’ (C1).
@@
Weight saliency mask is key to adapting MU for image generation. In Fig. 4, we explore the influence of the weight saliency mask in SalUn when transitioning to image generation. For comparison, we also include the random masking baseline. As we can see, the use of random masking in SalUn could lead to unstable generation performance when the condition is set to the forgetting class ‘airplane’ or other non-forgetting classes. First, the generation of noisy images may indicate over-forgetting, which contradicts the results obtained with Retrain as shown in Fig. 3. Second, when a non-forgetting class is specified, we notice the use of random masking degrades the generation quality, e.g., C2, C3, C6, and C7 in the figure. In contrast, our proposed SalUn, which leverages a proper weight saliency map, outperforms the implementation that uses random masking. This highlights the significance of proper weight saliency in MU for image generation. Extended from Fig. 4, we quantify the unlearning performance of Retrain, ESD, and SalUn using the two metrics introduced earlier, FID and UA, in Appendix B.3. With a comparable UA performance, it is worth highlighting that SalUn significantly outperforms ESD in FID. ESD seems to exhibit instability when forgetting and may generate low-quality images. Furthermore, we refer readers to Appendix B.1 for an ablation study on the sensitivity analysis of the weight saliency sparsity.
@@
Figure 5: Effectiveness of ‘nudity’ removal using different unlearned SD models acquired by SalUn, ESD, and FMN, respectively, and the original SD V1.4. The performance is measured by the # of generated harmful images against I2P prompts within each nudity category (i.e., row name). In the ‘Buttocks’ category, SalUn has no harmful generation.
Application to NSFW concept forgetting. Further, we assess the effectiveness of SalUn in concept-wise forgetting to eliminate the impact of NSFW (not safe for work) concepts introduced through inappropriate image prompts (I2P) (Schramowski et al., 2023). Specifically, we generate images from the open-source SD V1.4 using prompts provided by I2P and classify them into various corresponding nude body parts using the NudeNet detector (Bedapudi, 2019). Our goal is to leverage an MU method to effectively erase the influence of the nudity-related prompts in SD. Fig. 5 presents the unlearning performance of different unlearning methods, including SalUn, and the ESD and FMT baselines introduced in Table 2. Here, the unlearning effectiveness is measured by the amount of nudity-related image generations using the unlearned SD with I2P prompts. We also present the performance of the original SD model for comparison. As we can see, SalUn generates the fewest harmful images across all the nude body part classes. In particular, it significantly outperforms ESD (the second best-performing method) in ‘male breast’ and ‘female breast’. And the generated images no longer include exposed buttocks. Additionally, in the absence of unlearning, the original SD V1.4 can generate a large number of harmful images, underscoring the importance of MU for image generation. Fig. 6 provides a set of generated images using I2P prompts, illustrating the generation differences among the original SD model and its various unlearned versions.
@@
Figure 6: Examples of generated images using SDs w/ and w/o MU. The unlearning methods include ESD, FMN, and SalUn (ours). Each column contains generated images using different SDs with the same prompt (denoted by PPi) and the same seed. The detailed descriptions of the prompts used are provided in Table A9.